OpenStudy (anonymous):

I'm stuck: (x+y)^2/4x^2-y^2 =?? if x = 3, y = -5d. I got (3-5d)^2/26-(-5d)^2. I tried simplifying it even further but things got messy.

OpenStudy (anonymous):

is the eq (x+y)^2/(4x^2) -y^2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

x=3 and y =-5d (3+5d)^2/(4(3)^2)-(-5d)^2

OpenStudy (anonymous):

common denominator is 36

OpenStudy (anonymous):

so (3+5d)^2/36-(-5d)^2*36/36

OpenStudy (anonymous):

One question, shouldn't it be: (x+y)^2 = (3+(-5d))^2 = (3-5d)^2? Since a positive and a negative makes a negative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you are right.....

OpenStudy (anonymous):

but in this case there is a squared

OpenStudy (anonymous):

(x+y)^2=(3-5d)^2

OpenStudy (anonymous):

[(9-30d+25d^2)-(36*25d^2)]/36

OpenStudy (anonymous):

is that okay

OpenStudy (anonymous):

Okay, I'm getting it. I got 9-30d+25d^2/36+25d^2. Still trying to figure out the 36 part.

OpenStudy (anonymous):

Doesn't 25d^2 cancle each other out?

OpenStudy (anonymous):

no.... common denominator first

OpenStudy (anonymous):

Okay, one thing you answered with [(9-30d+25d^2)-(36*25d^2)]/36. Multiplication on the 2nd parenthesis not addition?

OpenStudy (anonymous):

you need to multiply both upper and lower fraction by the same number which in this case 36 for a common denominator